# Adaptive Random Testing with Qgrams: The Illusion Comes True

### v2

@@ 6.3. Research Questions and Metrics
The goal of our empirical study is to evaluate the test generation capabilities of ART with Qgram aggregation (Qgrams) in comparison to two competing techniques: ART with method sequence edit distance (Dist) and random testing (Rand).
RQ1 [Coverage] How does Qgrams compare to Dist and Rand in terms of coverage of the target web applications, under a fixed test generation time budget?
In this research question, we explore whether real-world web applications possess the characteristics that make Qgrams more effective than its competitors when given a fixed test generation budget (8 hours in our experiments).
@@
RQ2 [Efficiency] How does Qgrams compare to Dist and Rand in terms of the speed at which coverage grows as the number of generated tests increases?
@@
RQ3 [Uniqueness] To what extent does Qgrams cover unique test targets that the competing test generation techniques fail to cover?
This research question explores Qgrams’ ability to reach test targets that other approaches might miss. Since these unique targets could potentially contain faults, covering them suggests an increased likelihood of exposing bugs that other techniques might overlook.
@@
RQ4 [Sequence Length] What is the relationship between the length of the test sequence and its selection by Qgrams or Dist?
@@
Metrics. We measure the length of the test sequences selected by Qgrams or Dist over the test generation runs.
@@ 6.4. Experimental Procedure
In our empirical evaluation, we used the navigation graphs provided by Biagiola et al. (Biagiola et al., 2019) for the respective web applications, as well as their implemented POs. Given a navigation model for a certain web application, we follow the original paper and traverse the graph with a random walk of a length chosen randomly in the interval $[1,L]$, where $L=40$ is the maximum length. At each iteration, we randomly select an edge/method in the graph that is not yet covered, we generate a random walk starting from the index PO, and we connect the last node of the random walk with the target node of the uncovered edge. Considering Dist, Qgramss and Qgramss+i, we generate at each iteration $|W|=30$ candidates, and we execute the one that increases diversity the most according to the different methods (i.e., method sequence edit distance for Dist and entropy for Qgramss and Qgramss+i). For Qgramss and Qgramss+i, we use $Q=2$, i.e., we keep track of bigram counts during the execution.
We compared four techniques in our empirical study, i.e., Rand, Dist, Qgramss, and Qgramss+i, executing each of them with a time budget of eight hours. To account for the inherent randomness of the generation process, we executed each technique five times, and compared the results using rigorous statistical tests (Arcuri and Briand, 2014). In particular, we used the Wilcoxon rank sum test (Wilcoxon, 1945) to compute the $p$-value, with the confidence threshold $\alpha=0.05$, and the Vargha-Delaney effect size ($\hat{A}_{12}$(Vargha and Delaney, 2000) to assess the magnitude of the difference. Overall, we have six web applications, four techniques, an eight hour time budget, and five repetitions for a total CPU time of $6\times 4\times 8h\times 5=960h$ (or 40 CPU days).
@@ 6.5. Results
Table 4. Results for RQ1, RQ2, and RQ3. Web applications are ranked according to their complexity, shown in the first two columns. The grey-colored rows represent medium-complexity web applications, separating low-complexity (above), and high-complexity (below) applications. Bolded values represent the best average for each metric, while underlined values represent a $p$-value $<0.05$, and at least a medium effect size magnitude, of Qgrams w.r.t. Rand and Dist.
Cpx. Rand Dist Qgramss+i Qgramss
@@
Columns 3, 7, 11, and 15 of Table 4, show the average coverage, as a percentage of the total number of targets, achieved in five repetitions by Rand, Dist, Qgramss+i and Qgramss respectively, for each web application subject. We observe that in the web application subjects with low complexity, i.e., retroboard, and dimeshift, Rand has the highest coverage on average, although the difference with the other techniques is not statistically significant. On the other hand, for intermediate and high complexity web applications, Qgrams techniques outperform both Rand and Dist in most cases. Indeed, for phoenix, Qgramss covers significantly more than Rand and Dist with a large effect size (90% vs 75.79% and 73.16% respectively), while the difference with Qgramss+i is not statistically significant (Qgramss+i with an average coverage of 80% is significantly different from Dist with a large effect size). For splittypie, Qgramss+i covers significantly more than Rand and Dist with a large effect size (91.36% vs 82.73% and 82.27% respectively), while the difference with Qgramss (87.73%) is not statistically significant. For petclinic, Qgramss+i has the highest coverage, being very close to that of Qgramss (i.e., 83.40% vs 82.55%), although the difference with Rand and Dist is not statistically significant (both have an average coverage of 80%). For pagekit, both Qgramss+i and Qgramss+i cover significantly more than Rand and Dist, with a large effect size (81.89% and 83.59% vs 80.28% and 78.30%); although Qgramss has the highest average coverage, the difference with Qgramss+i is not statistically significant.
Figure 2. Coverage over number of executed tests for three subjects and all techniques   Rand,   Dist,   Qgramss,   Qgramss+i. Solid lines represent the average over five repetitions, while the shaded areas around them represent the standard error of the mean. Curves are padded with the respective last values, such that all techniques have curves with the same number of points. The left-hand side shows the coverage trend considering all executed test cases, with a zoomed-in view of the latest half of the executed test cases on the right. Best viewed in color.
@@
In retroboard, all techniques converge closely by the end, as shown by the overlapping uncertainties, with no technique clearly outperforming the others, though Rand (red line) achieves the highest average coverage. In phoenix, which contains harder-to-reach targets for Rand, the techniques are more spread out, with Qgramss (blue line) further apart, and Qgramss+i (orange line) slightly overlapping with Rand. In the high-complexity subject pagekit, the separation between Qgrams and the other techniques becomes even more pronounced, with both Qgramss and Qgramss+i showing a clear advantage over Rand and Dist, with no overlap.
RQ1 [Coverage]: In intermediate and high complexity web applications, Qgrams techniques outperform both Rand and Dist. In three out of four such cases, this improvement is statistically significant, with coverage gains ranging from 4% (petclinic) to 18% (splittypie).
RQ2 [Efficiency]. Columns 4, 8, 12, and 16 of Table 4, show the average area under the curve (AUC) over five repetitions for all the techniques and subjects. Areas have been normalized between 0 and 1, by dividing them by the maximum area (the area of the rectangle with base the maximum number of executed tests and height 100). The trend that we observe is similar to the coverage results. In low complexity web application subjects, Dist has the highest AUC (0.88 and 0.85 in retroboard and dimeshift respectively); Dist is significantly better than Rand (0.84) and Qgramss (0.85) with a large effect size in retroboard, and it is significantly better than Qgramss (0.75) with a large effect size in dimeshift. In phoenix and pagekit both Qgramss+i and Qgramss are significantly better than Rand and Dist with a large effect size. Qgramss has the highest average in both subjects (i.e., 0.84 and 0.80 respectively); the difference with Qgramss+i (0.77 and 0.79 respectively) is statistically significant only in pagekit. In splittypie, Qgramss+i (0.87) is significantly better than Rand (0.77), Dist (0.80) and Qgramss (0.82), while in petclinic both Qgramss+i and Qgramss achieve the best average AUC (i.e., 0.82), although only the difference with Rand (0.75) is statistically significant with a large effect size (the difference with Dist, with an AUC of 0.79, is not).
Columns 5, 9, 13, and 17 of Table 4, show the average area under the curve when considering 20% of the executed tests (AUC@20%). In this context, the differences are less pronounced than when considering AUC (only in phoenix, Qgramss is significantly better than Rand and Dist), although all diversity techniques significantly outperform Rand in five out of six applications (except dimeshift, and in pagekit Dist is not statistically different from Rand). Interestingly, Qgramss+i seems to always have the best AUC@20% average, that in low complexity subjects is matched by Dist, while in intermediate to high complexity is matched by Qgramss (except in splittypie, where Qgramss+i has the best average of all).
@@
Rand Dist Qgramss+i Qgramss
@@
Table 5 shows four statistics related to the test generation process for each technique and each subject. Particularly useful to interpret the efficiency results are the number of executed tests (Columns 1, 5, 9, and 13) and the time to generate a test case (Columns 4, 8, 12, and 16). The generation time for diversity-based techniques consists of the time to generate the $|W|=30$ candidate test cases (see Algorithm 1 and Algorithm 2), and the diversity maximization computation to select the test case that will be executed. On the other hand, the generation time for Rand only takes into account the time to generate a single test case. On average, considering all web application subjects (last row of Table 5), Rand only spends 0.02 $s$ in generating the test to execute, while Dist is much more expensive, taking on average 16.7 $s$ to generate a test case, as it scales quadratically with the number of executed test cases. On the other hand, Qgrams techniques are very efficient when selecting a test case to execute, taking on average 0.58 $s$ and 0.51 $s$ for Qgramss+i and Qgramss respectively, as they scale linearly with the number of executed tests. The cost for generating and selecting a test reflects on the total number of executed tests; indeed, Dist executes much less tests than Rand on average (i.e., 1,162 vs 3,877), while Qgrams techniques execute a comparable number of tests (3,894 and 3,807 respectively for Qgramss+i and Qgramss).
@@
The different number of tests executed by each technique is visible in Figure 2, where right-padding indicates that the budget finished (the curve was extended for the purpose of comparison with the other techniques). Dist spends most of the budget for diversity computation, with its coverage curves flattening much earlier than the competing techniques. On the other hand, Rand, by executing many more tests than Dist, is able to achieve a higher overall coverage than Dist if given enough time, although it is less efficient when a low test generation budget is available. Qgrams techniques combine the best of both worlds, i.e., the efficiency of Dist at low budget regimes (AUC@20%), and the ability to execute a comparable number of tests as Rand, at high budget regimes. This combination allows Qgrams techniques to be more efficient overall (AUC), especially in intermediate and high complexity subjects.
@@
RQ2 [Efficiency]: Qgrams techniques are significantly more efficient than Rand in five out of six subjects when the test generation budget is low (AUC@20%). With a higher test generation budget, Qgrams techniques outperform both Rand and Dist in intermediate and high complexity subjects, achieving statistical significance in three out of four cases, with AUC improvements ranging from 3.7% (petclinic) to 16.7% (phoenix).
@@
RQ3 [Uniqueness]. Columns 6, 10, 14 and 18 of Table 4, show the average number of coverage targets that a certain technique uniquely covers in each repetition. For instance, Rand covers on average 0.60 coverage targets in retroboard, that Dist, Qgramss+i and Qgramss do not cover; Rand has the highest average of uniquely covered targets in retroboard, but the difference with the other techniques is not statistically significant. In dimeshift, Qgramss+i has the highest average (1.40), above Qgramss (1.20), Rand (1.0) and Dist (0.80); again the difference with the other techniques is not significant. In phoenix, Qgramss covers significantly more unique targets (3.80) than all the other techniques with a large effect size, with Qgramss+i ranking second (0.60). Similarly, Qgramss covers significantly more unique targets (3.60) than the competing techniques with a large effect size in pagekit, with again Qgramss+i ranking second (1.40). On the other hand, Qgramss+i performs better than Qgramss in splittypie and petlicnic (1.60 and 1.20 vs 0.20 and 1.0 respectively); the difference with the other techniques is significant only in splittypie.
@@
RQ3 [Uniqueness]: Qgrams techniques outperform Rand and Dist in intermediate and high complexity web applications, achieving statistical significance in three out of four cases. In the worst case (pagekit), Qgrams techniques cover 4.5×4.5$\times$4.5 × more unique targets than the best of Rand and Dist, while in the best case (phoenix), they cover 19×19$\times$19 × more unique targets.
@@
RQ4 [Sequence Length]. Columns 2, 6, 10, and 14 of Table 5 show the average length (measured as number of statements) of the test cases selected for execution by each technique. We can observe that Dist selects the longest test cases, with an average length of 64.73 statements. On the other hand, Rand, Qgramss and Qgramss+i are quite comparable, featuring an average length of 39.04, 34.75, and 36.88 statements respectively. The length of the selected test case correlates with the time it takes to execute it, shown in Columns 3, 7, 11, and 15 of Table 5. Dist has the highest average with 5.04 $s$, followed by Rand with 4.19 $s$, Qgramss with 3.97 $s$ and Qgramss+i with 3.54 $s$. The length of the selected test case is the reason why Qgrams techniques are able to execute a comparable number of tests w.r.t. Rand, despite the higher generation time. In fact, the tests selected by Qgrams techniques are overall shorter than the tests generated by Rand (the only exception being petclinic, where both Qgramss+i and Qgramss select longer tests, and pagekit, where both select slightly longer tests), compensating the longer generation time.
@@
Figure 3. Length of the selected tests over the number of executed tests for three subjects and all techniques   Rand,   Dist,   Qgramss,   Qgramss+i. The length (y-axis) is expressed as number of statements. Solid lines represent the average over five repetitions, while the shaded areas around them represent the standard error of the mean. Each point is smoothed with a window size of 100, such that the trend is more visible. Curves are truncated at the minimum number of executed tests across techniques, to display the same number of points for all techniques. Best viewed in color.
@@
Figure 3 shows the trends (smoothed over a window of size 100) of the selected test case length over the number of executed tests of each technique for three subjects of different complexity, i.e., retroboard, phoenix and pagekit (full results are available in the replication package (testingautomated usi, 2024)). In all the plots, we observe that Dist tends to select the longest test cases to maximize diversity, and this does not change as the test generation progresses. This is natural given that a longer sequence allows more room for a high edit distance. The length of the test cases generated by Rand is lower, but it also remains roughly constant as the number of executed tests increases. On the other hand, both Qgrams techniques tend to select long tests at the beginning of the generation process, while the length decreases as executed tests accumulate. Indeed, maximization of the sequence edit distance can be achieved by selecting long test cases, while maximizing the entropy of the Qgrams requires selecting unexplored, but not necessarily long, sequences of statements. While selecting long sequences is beneficial to Qgrams at the beginning of the generation process, to maximize diversity and to quickly cover high probability targets, hard to cover targets need a more focused selection, which is facilitated by shorter, unexplored, sequences.
@@
RQ4 [Sequence Length]: Dist selects long test cases from the beginning to the end of the test generation process. On the other hand, Qgrams techniques tend to select long tests at the beginning, while privileging shorter tests at the end, focusing the generation budget towards unexplored sequences.

### v3

@@ 6.3. Research Questions and Metrics
The goal of our empirical study is to evaluate the test generation capabilities of ART with q-gram aggregation (q-grams) in comparison to two competing techniques: ART with method sequence edit distance (Dist) and random testing (Rand).
RQ1 [Coverage] How does q-grams compare to Dist and Rand in terms of coverage of the target web applications, under a fixed test generation time budget?
In this research question, we explore whether real-world web applications possess the characteristics that make q-grams more effective than its competitors when given a fixed test generation budget (8 hours in our experiments).
@@
RQ2 [Efficiency] How does q-grams compare to Dist and Rand in terms of the speed at which coverage grows as the number of generated tests increases?
@@
RQ3 [Uniqueness] To what extent does q-grams cover unique test targets that the competing test generation techniques fail to cover?
This research question explores q-grams’ ability to reach test targets that other approaches might miss. Since these unique targets could potentially contain faults, covering them suggests an increased likelihood of exposing bugs that other techniques might overlook.
@@
RQ4 [Sequence Length] What is the relationship between the length of the test sequence and its selection by q-grams or Dist?
@@
Metrics. We measure the length of the test sequences selected by q-grams or Dist over the test generation runs.
@@ 6.4. Experimental Procedure
In our empirical evaluation, we used the navigation graphs provided by Biagiola et al. (Biagiola et al., 2019) for the respective web applications, as well as their implemented POs. Given a navigation model for a certain web application, we follow the original paper and traverse the graph with a random walk of a length chosen randomly in the interval $[1,L]$, where $L=40$ is the maximum length. At each iteration, we randomly select an edge/method in the graph that is not yet covered, we generate a random walk starting from the index PO, and we connect the last node of the random walk with the target node of the uncovered edge. Considering Dist, q-gramss and q-gramss+i, we generate at each iteration $|W|=30$ candidates, and we execute the one that increases diversity the most according to the different methods (i.e., method sequence edit distance for Dist and entropy for q-gramss and q-gramss+i). For q-gramss and q-gramss+i, we use $Q=2$, i.e., we keep track of bigram counts during the execution.
We compared four techniques in our empirical study, i.e., Rand, Dist, q-gramss, and q-gramss+i, executing each of them with a time budget of eight hours. To account for the inherent randomness of the generation process, we executed each technique five times, and compared the results using rigorous statistical tests (Arcuri and Briand, 2014). In particular, we used the Wilcoxon rank sum test (Wilcoxon, 1945) to compute the $p$-value, with the confidence threshold $\alpha=0.05$, and the Vargha-Delaney effect size ($\hat{A}_{12}$(Vargha and Delaney, 2000) to assess the magnitude of the difference. Overall, we have six web applications, four techniques, an eight hour time budget, and five repetitions for a total CPU time of $6\times 4\times 8h\times 5=960h$ (or 40 CPU days).
@@ 6.5. Results
Table 4. Results for RQ1, RQ2, and RQ3. Web applications are ranked according to their complexity, shown in the first two columns. The grey-colored rows represent medium-complexity web applications, separating low-complexity (above), and high-complexity (below) applications. Bolded values represent the best average for each metric, while underlined values represent a $p$-value $<0.05$, and at least a medium effect size magnitude, of q-grams w.r.t. Rand and Dist. For each of the four evaluation metrics (AUC, AUC@20%, Unique Targets, Coverage) and for each web app, bolded values represent the best average among Rand, Dist, q-gramss+i, q-gramss. Sometimes the best average is a tie involving 2 or more techniques (e.g., pagekit has a tie on AUC@20, with the best average value of 0.14 reached by Dist, q-gramss+i, and q-gramss).
Cpx. Rand Dist q-gramss+i q-gramss
@@
Columns 3, 7, 11, and 15 of Table 4, show the average coverage, as a percentage of the total number of targets, achieved in five repetitions by Rand, Dist, q-gramss+i and q-gramss respectively, for each web application subject. We observe that in the web application subjects with low complexity, i.e., retroboard, and dimeshift, Rand has the highest coverage on average, although the difference with the other techniques is not statistically significant. On the other hand, for intermediate and high complexity web applications, q-grams techniques outperform both Rand and Dist in most cases. Indeed, for phoenix, q-gramss covers significantly more than Rand and Dist with a large effect size (90% vs 75.79% and 73.16% respectively), while the difference with q-gramss+i is not statistically significant (q-gramss+i with an average coverage of 80% is significantly different from Dist with a large effect size). For splittypie, q-gramss+i covers significantly more than Rand and Dist with a large effect size (91.36% vs 82.73% and 82.27% respectively), while the difference with q-gramss (87.73%) is not statistically significant. For petclinic, q-gramss+i has the highest coverage, being very close to that of q-gramss (i.e., 83.40% vs 82.55%), although the difference with Rand and Dist is not statistically significant (both have an average coverage of 80%). For pagekit, both q-gramss+i and q-gramss+i cover significantly more than Rand and Dist, with a large effect size (81.89% and 83.59% vs 80.28% and 78.30%); although q-gramss has the highest average coverage, the difference with q-gramss+i is not statistically significant.
Figure 2. Coverage over number of executed tests for three subjects and all techniques   Rand,   Dist,   q-gramss,   q-gramss+i. Solid lines represent the average over five repetitions, while the shaded areas around them represent the standard error of the mean. Curves are padded with the respective last values, such that all techniques have curves with the same number of points. The left-hand side shows the coverage trend considering all executed test cases, with a zoomed-in view of the latest half of the executed test cases on the right. Best viewed in color.
@@
In retroboard, all techniques converge closely by the end, as shown by the overlapping uncertainties, with no technique clearly outperforming the others, though Rand (red line) achieves the highest average coverage. In phoenix, which contains harder-to-reach targets for Rand, the techniques are more spread out, with q-gramss (blue line) further apart, and q-gramss+i (orange line) slightly overlapping with Rand. In the high-complexity subject pagekit, the separation between q-grams and the other techniques becomes even more pronounced, with both q-gramss and q-gramss+i showing a clear advantage over Rand and Dist, with no overlap.
RQ1 [Coverage]: In intermediate and high complexity web applications, q-grams techniques outperform both Rand and Dist. In three out of four such cases, this improvement is statistically significant, with coverage gains ranging from 4% (petclinic) to 18% (splittypie).
RQ2 [Efficiency]. Columns 4, 8, 12, and 16 of Table 4, show the average area under the curve (AUC) over five repetitions for all the techniques and subjects. Areas have been normalized between 0 and 1, by dividing them by the maximum area (the area of the rectangle with base the maximum number of executed tests and height 100). The trend that we observe is similar to the coverage results. In low complexity web application subjects, Dist has the highest AUC (0.88 and 0.85 in retroboard and dimeshift respectively); Dist is significantly better than Rand (0.84) and q-gramss (0.85) with a large effect size in retroboard, and it is significantly better than q-gramss (0.75) with a large effect size in dimeshift. In phoenix and pagekit both q-gramss+i and q-gramss are significantly better than Rand and Dist with a large effect size. q-gramss has the highest average in both subjects (i.e., 0.84 and 0.80 respectively); the difference with q-gramss+i (0.77 and 0.79 respectively) is statistically significant only in pagekit. In splittypie, q-gramss+i (0.87) is significantly better than Rand (0.77), Dist (0.80) and q-gramss (0.82), while in petclinic both q-gramss+i and q-gramss achieve the best average AUC (i.e., 0.82), although only the difference with Rand (0.75) is statistically significant with a large effect size (the difference with Dist, with an AUC of 0.79, is not).
Columns 5, 9, 13, and 17 of Table 4, show the average area under the curve when considering 20% of the executed tests (AUC@20%). In this context, the differences are less pronounced than when considering AUC (only in phoenix, q-gramss is significantly better than Rand and Dist), although all diversity techniques significantly outperform Rand in five out of six applications (except dimeshift, and in pagekit Dist is not statistically different from Rand). Interestingly, q-gramss+i seems to always have the best AUC@20% average, that in low complexity subjects is matched by Dist, while in intermediate to high complexity is matched by q-gramss (except in splittypie, where q-gramss+i has the best average of all).
@@
Rand Dist q-gramss+i q-gramss
@@
Table 5 shows four statistics related to the test generation process for each technique and each subject. Particularly useful to interpret the efficiency results are the number of executed tests (Columns 1, 5, 9, and 13) and the time to generate a test case (Columns 4, 8, 12, and 16). The generation time for diversity-based techniques consists of the time to generate the $|W|=30$ candidate test cases (see Algorithm 1 and Algorithm 2), and the diversity maximization computation to select the test case that will be executed. On the other hand, the generation time for Rand only takes into account the time to generate a single test case. On average, considering all web application subjects (last row of Table 5), Rand only spends 0.02 $s$ in generating the test to execute, while Dist is much more expensive, taking on average 16.7 $s$ to generate a test case, as it scales quadratically with the number of executed test cases. On the other hand, q-grams techniques are very efficient when selecting a test case to execute, taking on average 0.58 $s$ and 0.51 $s$ for q-gramss+i and q-gramss respectively, as they scale linearly with the number of executed tests. The cost for generating and selecting a test reflects on the total number of executed tests; indeed, Dist executes much less tests than Rand on average (i.e., 1,162 vs 3,877), while q-grams techniques execute a comparable number of tests (3,894 and 3,807 respectively for q-gramss+i and q-gramss).
@@
The different number of tests executed by each technique is visible in Figure 2, where right-padding indicates that the budget finished (the curve was extended for the purpose of comparison with the other techniques). Dist spends most of the budget for diversity computation, with its coverage curves flattening much earlier than the competing techniques. On the other hand, Rand, by executing many more tests than Dist, is able to achieve a higher overall coverage than Dist if given enough time, although it is less efficient when a low test generation budget is available. q-grams techniques combine the best of both worlds, i.e., the efficiency of Dist at low budget regimes (AUC@20%), and the ability to execute a comparable number of tests as Rand, at high budget regimes. This combination allows q-grams techniques to be more efficient overall (AUC), especially in intermediate and high complexity subjects.
@@
RQ2 [Efficiency]: q-grams techniques are significantly more efficient than Rand in five out of six subjects when the test generation budget is low (AUC@20%). With a higher test generation budget, q-grams techniques outperform both Rand and Dist in intermediate and high complexity subjects, achieving statistical significance in three out of four cases, with AUC improvements ranging from 3.7% (petclinic) to 16.7% (phoenix).
@@
RQ3 [Uniqueness]. Columns 6, 10, 14 and 18 of Table 4, show the average number of coverage targets that a certain technique uniquely covers in each repetition. For instance, Rand covers on average 0.60 coverage targets in retroboard, that Dist, q-gramss+i and q-gramss do not cover; Rand has the highest average of uniquely covered targets in retroboard, but the difference with the other techniques is not statistically significant. In dimeshift, q-gramss+i has the highest average (1.40), above q-gramss (1.20), Rand (1.0) and Dist (0.80); again the difference with the other techniques is not significant. In phoenix, q-gramss covers significantly more unique targets (3.80) than all the other techniques with a large effect size, with q-gramss+i ranking second (0.60). Similarly, q-gramss covers significantly more unique targets (3.60) than the competing techniques with a large effect size in pagekit, with again q-gramss+i ranking second (1.40). On the other hand, q-gramss+i performs better than q-gramss in splittypie and petlicnic (1.60 and 1.20 vs 0.20 and 1.0 respectively); the difference with the other techniques is significant only in splittypie.
@@
RQ3 [Uniqueness]: q-grams techniques outperform Rand and Dist in intermediate and high complexity web applications, achieving statistical significance in three out of four cases. In the worst case (pagekit), q-grams techniques cover 4.5×4.5$\times$4.5 × more unique targets than the best of Rand and Dist, while in the best case (phoenix), they cover 19×19$\times$19 × more unique targets.
@@
RQ4 [Sequence Length]. Columns 2, 6, 10, and 14 of Table 5 show the average length (measured as number of statements) of the test cases selected for execution by each technique. We can observe that Dist selects the longest test cases, with an average length of 64.73 statements. On the other hand, Rand, q-gramss and q-gramss+i are quite comparable, featuring an average length of 39.04, 34.75, and 36.88 statements respectively. The length of the selected test case correlates with the time it takes to execute it, shown in Columns 3, 7, 11, and 15 of Table 5. Dist has the highest average with 5.04 $s$, followed by Rand with 4.19 $s$, q-gramss with 3.97 $s$ and q-gramss+i with 3.54 $s$. The length of the selected test case is the reason why q-grams techniques are able to execute a comparable number of tests w.r.t. Rand, despite the higher generation time. In fact, the tests selected by q-grams techniques are overall shorter than the tests generated by Rand (the only exception being petclinic, where both q-gramss+i and q-gramss select longer tests, and pagekit, where both select slightly longer tests), compensating the longer generation time.
@@
Figure 3. Length of the selected tests over the number of executed tests for three subjects and all techniques   Rand,   Dist,   q-gramss,   q-gramss+i. The length (y-axis) is expressed as number of statements. Solid lines represent the average over five repetitions, while the shaded areas around them represent the standard error of the mean. Each point is smoothed with a window size of 100, such that the trend is more visible. Curves are truncated at the minimum number of executed tests across techniques, to display the same number of points for all techniques. Best viewed in color.
@@
Figure 3 shows the trends (smoothed over a window of size 100) of the selected test case length over the number of executed tests of each technique for three subjects of different complexity, i.e., retroboard, phoenix and pagekit (full results are available in the replication package (testingautomated usi, 2024)). In all the plots, we observe that Dist tends to select the longest test cases to maximize diversity, and this does not change as the test generation progresses. This is natural given that a longer sequence allows more room for a high edit distance. The length of the test cases generated by Rand is lower, but it also remains roughly constant as the number of executed tests increases. On the other hand, both q-grams techniques tend to select long tests at the beginning of the generation process, while the length decreases as executed tests accumulate. Indeed, maximization of the sequence edit distance can be achieved by selecting long test cases, while maximizing the entropy of the q-grams requires selecting unexplored, but not necessarily long, sequences of statements. While selecting long sequences is beneficial to q-grams at the beginning of the generation process, to maximize diversity and to quickly cover high probability targets, hard to cover targets need a more focused selection, which is facilitated by shorter, unexplored, sequences.
@@
RQ4 [Sequence Length]: Dist selects long test cases from the beginning to the end of the test generation process. On the other hand, q-grams techniques tend to select long tests at the beginning, while privileging shorter tests at the end, focusing the generation budget towards unexplored sequences.